OpenStudy (winsteria):

A recipe that makes 7 servings calls for 1/2 cup of apple juice and 2/3 cup of cranberry juice. How many cups of juice are there per serving? A. 1/6 B. 3/5 C. 1(and) 1/16 D. 8 (and) 1/6 PLEASE HELP???

OpenStudy (winsteria):

PLEASEEE

OpenStudy (asevilla5):

im working on it one sec @winsteria

OpenStudy (winsteria):

okay thank you so much take your time

OpenStudy (asevilla5):

@Hero

OpenStudy (asevilla5):

hard time getting the answer. here is a link to a similar problem like this http://openstudy.com/study#/updates/53f22ac4e4b01789aba7847b

OpenStudy (anonymous):

First add your fractions to determine the total juice

OpenStudy (anonymous):

1/2 + 2/3 = 3/6 + 4/6 = 7/6 = 1 and 1/6

OpenStudy (anonymous):